A regression line was calculated as y = 9.7 - 3.2x. What is the slope of this line?
The equation of a regression line is y=4.6 +3.2x.Find the y intercept of this line.
If the correlation coefficient r is equal to .53, find the coefficient of determination and the coefficient f nondetermination.

Answer by stanbon(75887)

A regression line was calculated as y = 9.7 - 3.2x. What is the slope of this line?
slope = -3.2
---------------------------
The equation of a regression line is y=4.6 +3.2x.Find the y intercept of this line.
intercept = 4.6
--------------------------
If the correlation coefficient r is equal to .53, find the coefficient of determination
R^2 = 0.2809
------------------------------
and the coefficient f nondetermination
1-R^2 = 0.7191
============================================
Cheers,
Stan H.